OpenStudy (anonymous):

idk if you know this one The probability that a train leaves on time is 0.9. The probability that the train arrives on time and leaves on time is 0.36. What is the probability that the train arrives on time given that it leaves on time? 0.6 0.9 0.27 0.36

OpenStudy (anonymous):

wow this is a crappy problem quite the collection of topics you have tonight isn't it?

OpenStudy (anonymous):

i will tell you what they want you to do then want you to compute \[\frac{.36}{.9}\]

jimthompson5910 (jim_thompson5910):

A = arrives on time L = leaves on time we're given P(L) = 0.9 and P(A and L) = 0.36 They want you to find P(A|L) and you would use the conditional probability formula P(A|L) = P(A and L)/P(L) to compute

OpenStudy (anonymous):

i got 0.4 0.o @satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@satellite73 but thats not a answer choice ?

OpenStudy (anonymous):

really? what are the choices?

OpenStudy (anonymous):

0.6 0.9 0.27 0.36

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

sorry i have to stick with my answer

OpenStudy (anonymous):

i have two questions left if you could help me @satellite73

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

the total number in state A is 40 the number of liberals is 22 your answer is \(\frac{22}{40}=\frac{11}{20}\)

OpenStudy (anonymous):

or if you prefer, \(55\%\)

OpenStudy (anonymous):

ok now i want to know, before we do the last one, how you got such a random selection of problems all different topics now trig! and not even a basic trig question, a harder one

OpenStudy (anonymous):

my teacher @satellite73

OpenStudy (anonymous):

its a review @satellite73

OpenStudy (anonymous):

a review of what? fifty different non related topics? the answer to the last one is D, but my guess is you do not know why am i right?

OpenStudy (anonymous):

no lol but thank you for all your help:) @satellite73